OpenStudy (8bitfox):

Can someone help me solve inequalities |5-m| >or equal to 7 Can someone help explain this to me??

OpenStudy (blanablas):

I can help explain

OpenStudy (blanablas):

You understand absolute value and inequalities, right?

OpenStudy (8bitfox):

yes for example |-2| =2 |3| = 3 no matter what its sign is right?

OpenStudy (blanablas):

yes

OpenStudy (blanablas):

so for every value in there, it must always be greater than 7

OpenStudy (blanablas):

are you graphing this?

OpenStudy (8bitfox):

what do you mean by graphing?

OpenStudy (blanablas):

Did they ask you to graph it, or is this the entire question

OpenStudy (blanablas):

inequalites can be graphed

OpenStudy (blanablas):

but i'll go ahead and show u how to solve it

OpenStudy (blanablas):

so absolute values don't have a real opposite, like addition and subtraction for example

OpenStudy (blanablas):

so you just ignore them really

OpenStudy (8bitfox):

Yeah they do

OpenStudy (blanablas):

they do? i think i'm having a brain fart here

OpenStudy (8bitfox):

one second

OpenStudy (blanablas):

ok

OpenStudy (blanablas):

Actually, thinking about this equation

OpenStudy (blanablas):

every value for x < -2 and x > 12 works

OpenStudy (blanablas):

so there's a break in the line, and I never remember doing that

OpenStudy (8bitfox):

They ask me to state the equation or inequality as a disjunction or conjuction then solve it and graph it

OpenStudy (blanablas):

OH i remember

OpenStudy (blanablas):

its a disjunction

OpenStudy (blanablas):

to solve it, you can use the +/- sign

OpenStudy (blanablas):

solve for m - 5 and -m + 5

OpenStudy (blanablas):

basically flipping the signs in the second one

OpenStudy (blanablas):

solving for m - 5, it's m >/= 12 solving for -m + 5, it's -m < 2 m = -2

OpenStudy (blanablas):

made a typo, i meant m>/=-2

OpenStudy (blanablas):

and when you flip the signs, always turn the inequality sign around

OpenStudy (blanablas):

So to graph it

OpenStudy (blanablas):

there are two lines drawn

OpenStudy (blanablas):

one for -m + 5 < 7 and one for m - 5 > 7

OpenStudy (blanablas):

the line will be a solid one, not dashed as it's a </= sign

OpenStudy (8bitfox):

okay give me a second im trying to grasp all of this

OpenStudy (blanablas):

do i need to explain the graph?

OpenStudy (blanablas):

Ok, ask if you need help

OpenStudy (blanablas):

brb, need to submit a test i'm taking

OpenStudy (blanablas):

few seconds

OpenStudy (8bitfox):

so for example FI have |5-m| >or equal to 7 FIRST I change that to |5-m| >/ 7 AND |5-m| \< 7 correct?

OpenStudy (8bitfox):

Then wouldn't I have to add 5 to both sides??

OpenStudy (blanablas):

back

OpenStudy (8bitfox):

oops wait hold on I did that for the wrong inequality thats for less than

OpenStudy (blanablas):

Ok, so m-5>= 7, and -m+5 =<7

OpenStudy (blanablas):

Without it, it would be a regular inequality

OpenStudy (blanablas):

I'll list the steps for solving both equations

OpenStudy (blanablas):

For |m-5| >= 7: |m| >= 12, but it should be written as m >= 12

OpenStudy (blanablas):

what i did there is add 5 to both sides

OpenStudy (blanablas):

For |-m+5| < 7: |-m| < 2 (-5 both sides) |m| < -2 (multiplied by -1)

OpenStudy (blanablas):

i'm just going to delete the spam omments and explain it all in one mega-comment

OpenStudy (8bitfox):

im so confused @~@

OpenStudy (blanablas):

yeah i'm sorry

OpenStudy (blanablas):

Ok here goes

OpenStudy (blanablas):

To solve an inequality with an absolute value sign, you must solve the original and inverted form of it

OpenStudy (blanablas):

The original, in this case, is |m-5| >= 7

OpenStudy (blanablas):

the inverted is multiplying the whole left side by -1

OpenStudy (blanablas):

so |-m+5| <= 7 (and flipping the inequality sign)

OpenStudy (blanablas):

Then, to make things simple, just take away the absolute value signs and solve for m

OpenStudy (8bitfox):

first why do yo you multiply it by -1 (sorry im just trying me best to undertand this step by step

OpenStudy (blanablas):

because absolute values don't care about negatives

OpenStudy (8bitfox):

alright

OpenStudy (blanablas):

if m was -2, it would solve to |-2 -5| = 7 |-7| = 7

OpenStudy (blanablas):

that's why

OpenStudy (blanablas):

so solving the original, we get: |m-5| >=7 |m| >= 12 (add 5 both sides)

OpenStudy (blanablas):

and the inverted: -m + 5 <= 7 -m=2 (subtract 5 both sides) m=2(multiply by -1)

OpenStudy (blanablas):

Now for graphing it

OpenStudy (blanablas):

Because m >= 12 OR m<=-2, it is a disjunction

OpenStudy (blanablas):

or = disjunction

OpenStudy (blanablas):

What you do here is you remove the inequality sign for both the original and inverted equation, and replace them with a = sign

OpenStudy (blanablas):

so you can have the lines for your graph there are going to be two lines, you shade in a certain part of them

OpenStudy (blanablas):

wait a minute, this is a 1-variable equation

OpenStudy (blanablas):

you can only plot it on a number line

OpenStudy (blanablas):

if they say to graph it, then somethings wrong

OpenStudy (blanablas):

to graph this, just put points at -2 and 12 on a number line (make sure they are filled in dots because of the >= sign) and draw lines going outwards

OpenStudy (blanablas):

and that's it

OpenStudy (blanablas):

sorry for confusing you at first

OpenStudy (8bitfox):

let me check the paper again

OpenStudy (blanablas):

ok

OpenStudy (8bitfox):

and don't worry about it thats fine

OpenStudy (blanablas):

if they say to graph it, they're wrong, just draw a number line and inform whoever gave it to you of their mistake

OpenStudy (blanablas):

unless by graph they mean graph a number line which sounds weird

OpenStudy (8bitfox):

they don't really explain it they just say to graph it, don't worry about it I'll talk to them and explain but thank you anyway for helping with this question

OpenStudy (blanablas):

ok, np sorry for confusing you at first

OpenStudy (blanablas):

tag me if you need help

OpenStudy (8bitfox):

I will, again thank you :)